# Co-Writing with AI, on Human Terms: Aligning Research with User Demands Across the Writing Process

### v1

@@ 6.1. Key Findings
6.1.1. RQ 1: What design strategies are used or suggested in existing AI-assisted writing research and how are these strategies distributed across writing processes and contexts?
We answered the first research questions through our systematic review and thematic analysis (section 3), where we identified four primary strategies for AI-assisted writing support: Structured Guidance (S1), Guided Exploration (S2), Active Co-Writing (S3), and Critical Feedback (S4). S1 provides structured guidance while building user skills (e.g., LitWeaver by Choe et al. (Choe et al., 2024) leads novice researchers through completing a literature review), S2 enables creative control through systematic exploration (e.g., ABScribe by Reza et al. (Reza et al., 2024) enables users to rapidly iterate on chunks of text, storing previously-explored ideas and recipes for future exploration and revision), S3 supports efficient collaboration while maintaining user control (e.g. DiaryMate by Kim et al. (Kim et al., 2024b) encourages users to select between AI suggestions to compose a diary that was meaningful to them), and S4 promotes strategies that facilitate user reflection and engagement through analysis and feedback (e.g. Impressona by Benharrak et al. (Benharrak et al., 2024) specifies Personas that provide targeted feedback, prompting user reflection with a particular audience in mind). These strategies are valuable because they distill Generative AI research into actionable insights from the literature.
@@
6.1.2. RQ2: Which cognitive processes do writers consider essential to control in order to maintain their sense of agency during AI-assisted writing, and how do user situations, writing contexts, and AI interaction types shape their perceptions of ownership?
While the strategies represent current research, they do not offer guidance on writers’ values tied to preserving human agency. Our second study helps bridge this gap. We found three themes that explain when ownership matters (in relation to four contextual factors: time, importance, confidence, and trust, covered in Section 4.2.1), what writers want to own (in relation to two primary contribution types: content and form, covered in Section Section 4.2.2), across the cognitive processes: planning, reviewing, and translating (Flower and Hayes, 1981).)
@@
Each study offers useful insights on their own, but combining them is far more useful to designers because together, it not only maps the current research landscape, but also enables us to offer designers guidance on what should be done to align with user demands, as explored in detail in section 5. Study 1 is akin to a map handed to a sailor (the designer). Study 2 is akin to a compass that tells them where to go. Our findings indicate how writers’ sense of ownership is tied to specific cognitive processes: content-focused writers derive ownership primarily from ideation during the planning phase, as they feel that is where their primary contribution lies. In contrast, form-oriented writers connect their sense of ownership to translation and review, as that is where they want to exercise control over stylistic elements. This view of ownership suggests a ‘chessboard-like’ pattern, where users seek AI assistance in areas outside their primary contribution. Table 3 illustrates this preference: green regions show where AI support is sought, red areas denote places where AI should not intervene, and yellow regions denote zones where AI may assist with caution.
@@ 6.2. Contributions to CSCW
This example illustrates the value of pairing systematic reviews with user studies that go deep into areas of interest and importance to the research community, such as our focus on preserving human agency. Within that context, our work relates to broader theories in Human-Centered AI, such as Ben Shneiderman’s HAI framework, which argues that automation and human control need not be at odds on a unidimensional spectrum (Shneiderman, 2022), like in the classic 1978 characterization of automation by Sheridan and Verplank (Sheridan et al., 1978). Instead, Shneiderman posits a multidimensional perspective where automation and control can increase concurrently, which resonates with our optimistic vision for AI’s role in augmenting human agency. Like Shneiderman’s multi-dimensional characterization of automation and human agency, our approach demonstrates the value of viewing creative tasks as a multi-dimensional. By breaking it down into distinct cognitive processes and contexts, we move beyond a one-size-fits-all perspective and highlight specific context $\times$× process dimensions where designers should focus AI support.
@@ 6.3. Limitations and Future Work
Thirdly, while our inclusion criteria was broad, allowing participants aged 18 and above, the second requirement that participants have some prior experience using AI tools for writing inadvertently limited the age diversity in our sample, resulting in a maximum age of 34. This excludes valuable insights from older adults who are also impacted by AI. Future studies could address this by incorporating a more representative age distribution to explore potential age-based differences in attitudes toward AI-assisted writing and ownership. The gender composition of our sample could be expanded to examine gender-specific perspectives. Furthermore, as our study only included participants familiar with AI, our findings are less applicable to writers with no prior familiarity. Future research could investigate the initial reactions and adoption experiences of AI-naive writers, illuminating potential barriers to entry and differing perceptions of agency in AI-assisted writing.
@@ 7. Conclusion
Acknowledgements.
The authors acknowledge the use of ChatGPT, a generative AI tool developed by OpenAI, in the writing process of this paper. The tool was used solely to enhance the readability of specific sections. The authors retain responsibility and ownership of all core ideas, concepts, and interpretations. This usage aligns with ACM policy, which requires authors to be the ”creator or originator of an idea.”

### v2

@@ 6.1. Key Findings
To structure our discussion of key findings, we divide the two research questions into four analytical components:
  • RQ1.1 – Design Strategies: What design strategies are currently used or proposed in AI-assisted writing research, particularly in terms of interaction models and intended uses of AI-generated output?
@@
  • RQ1.2 – Process & Context Distribution: How are these design strategies distributed across different cognitive processes and writing contexts?
@@
  • RQ2.1 – Agency-Preserving Processes: Which cognitive processes do writers consider essential to control in order to maintain their sense of agency during AI-assisted writing?
  • RQ2.2 – Modulating Factors in Ownership: How do user situations, writing contexts, and AI interaction types shape writers’ perceptions of ownership?
We answered the first research question encompassing components 1.1 and 1.2 through our systematic review and thematic analysis (Section 3).
6.1.1. RQ 1.1 Design Strategies
We identified four primary strategies for AI-assisted writing support that embody the existing literature: Structured Guidance (S1), Guided Exploration (S2), Active Co-Writing (S3), and Critical Feedback (S4):
  • S1: Structured Guidance supports users by providing step-by-step AI-assistance while fostering skill development (e.g., LitWeaver by Choe et al. (Choe et al., 2024) leads novice researchers through completing a literature review)
  • S2: Guided Exploration enables creative control through systematic exploration (e.g., ABScribe by Reza et al. (Reza et al., 2024) enables users to rapidly iterate on chunks of text, storing previously-explored ideas and recipes for future exploration and revision)
  • S3 Active Co-Writing facilitates efficient collaboration while maintaining user control (e.g. DiaryMate by Kim et al. (Kim et al., 2024b) encourages users to select between AI suggestions to compose a diary that was meaningful to them)
  • S4: Critical Feedback promotes strategies that facilitate user reflection and engagement through analysis and feedback (e.g. Impressona by Benharrak et al. (Benharrak et al., 2024) specifies Personas that provide targeted feedback, prompting user reflection with a particular audience in mind)
6.1.2. RQ 1.2 Process & Context Distribution
To identify how the design strategies are distributed across different cognitive processes and writing contexts, we mapped our dataset across five writing contexts adapted from Lee et al. (Lee et al., 2024)—Academic, Creative, Formal, Personal, and General—and three cognitive processes based on Flower and Hayes (Hayes, 1996)—Planning, Translating, Reviewing. This mapping revealed uneven research attention across processes and contexts (shown in Table 1), such as strong representation of AI support during the Generating subprocess within the Planning phase in Creative and Academic writing contexts, but limited focus on Monitoring support across all writing contexts.
Our investigation also revealed how these strategies are distributed across writing processes and contexts (see Table 2), with S3 being the most widely used strategy overall (37.1% of all systems in the dataset). Most notably, Creative writing showed a surprisingly high deployment of S3 systems (50%), highlighting a potential misalignment between strategy use and user needs, as Active Co-Writing (S3) can compromise the sense of ownership central to writers’ creative expression. We also found comparatively lower deployment of S4 systems (Critical Feedback), which accounted for only 14.5% of the dataset. This uneven distribution suggests opportunities for expanding underused strategies like S4, particularly in writing contexts and processes where reflection and iteration play a central role.
While the strategies represent current research, they do not offer guidance on writers’ values tied to preserving human agency. Our second study helps bridge this gap, and helps answer the second research question comprising components 2.1 and 2.2.
6.1.3. RQ 2.1 Agency-Preserving Processes:
Our interview findings show that writers consider different cognitive processes as essential depending on what they see as their core creative contribution. Planning is deemed critical by content-oriented writers (e.g., academics and technical writers). These writers derive their ownership from generating original ideas, and prioritize information structure over style. Therefore, delegating this process to AI would undermine their sense of ownership. On the other hand, Translation and Revision are deemed essential by form focused-writers (e.g., poets and novelists), as their ownership is rooted in their voice, tone, and sentence or word-level decisions. Maintaining control over these granular stylistic dimensions is central to preserving their identity as writers.
6.1.4. RQ 2.2 Modulating Factors in Ownership:
Writers’ perceptions of ownership are flexible, and are modulated by several factors, which we organize into three major themes spanning different user situations, which influence when ownership matters, writing contexts influencing what writers want to own , and AI interaction types, which affect how writers experience agency during AI-assisted writing. The first theme, covered in Section 4.2.1, includes four key situational factors: time, importance, confidence, and trust. The second theme, covered in Section 4.2.2, distinguishes between the two primary contribution types: content and form. The third theme, covered in Section 4.2.3, highlights four interaction types—suggestions, final say, global AI toggle, and local AI toggle—that modulate writers’ sense of control when co-writing with AI. Together, these factors reveal that ownership in AI-assisted writing is not fixed but actively negotiated—shaped by writers’ goals, contexts, and the interaction design of the tools they use.
@@
Each study offers useful insights on its own, but in combination they are more useful to designers because together they not only map the current research landscape, but also enable us to offer designers guidance on what should be done to align with user demands, as explored in detail in Section 5. Study 1 is akin to a map handed to a sailor (the designer). Study 2 is akin to a compass that tells them where to go. Our findings indicate how writers’ sense of ownership is tied to specific cognitive processes: content-focused writers derive ownership primarily from ideation during the planning phase, as they feel that is where their primary contribution lies. In contrast, form-oriented writers connect their sense of ownership to translation and review, as that is where they want to exercise control over stylistic elements. This view of ownership suggests a ‘chessboard-like’ pattern, where users seek AI assistance in areas outside their primary contribution. Table 3 illustrates this preference: blue regions show where AI support is sought, pink areas denote places where AI should not intervene, and yellow regions denote zones where AI may assist with caution.
@@ 6.2. Contributions to CSCW
This example illustrates the value of pairing systematic reviews with user studies that go deep into areas of interest and importance to the research community, such as our focus on preserving human agency. Within that context, our work relates to broader theories in Human-Centered AI, such as Ben Shneiderman’s HAI framework, which argues that automation and human control need not be at odds on a unidimensional spectrum (Shneiderman, 2022), like in the classic 1978 characterization of automation by Sheridan and Verplank (Sheridan et al., 1978). Instead, Shneiderman posits a multidimensional perspective where automation and control can increase concurrently, which resonates with our optimistic vision for AI’s role in augmenting human agency. Like Shneiderman’s multi-dimensional characterization of automation and human agency, our approach demonstrates the value of viewing creative tasks as a multi-dimensional. By breaking it down into distinct cognitive processes and contexts, we move beyond a one-size-fits-all perspective and highlight specific context $\times$ process dimensions where designers should focus AI support.
@@ 6.3. Limitations and Future Work
Thirdly, while our inclusion criteria were broad, allowing participants aged 18 and above, the second requirement that participants have some prior experience using AI tools for writing inadvertently limited the age diversity in our sample, resulting in a maximum age of 34. This excludes valuable insights from older adults who are also impacted by AI. Additionally, our sample exhibited limited educational diversity, with all but one participant having attended university. This educational homogeneity may have influenced perspectives on AI-assisted writing, as individuals with higher educational backgrounds might approach AI tools differently than those with other experiences. Given that LLMs serve as writing tools for many people beyond academic contexts, future studies should prioritize more educationally- and age-diverse samples to capture how different populations perceive and interact with AI writing assistance and ownership. The gender composition of our sample could also be expanded to examine gender-specific perspectives. Furthermore, as our study only included participants familiar with AI, our findings are less applicable to writers with no prior familiarity. Future research could investigate the initial reactions and adoption experiences of AI-naive writers, illuminating potential barriers to entry and differing perceptions of agency in AI-assisted writing.